# Driving as a Diagnostic Tool: Scenario-based Cognitive Assessment in Older Drivers from Driving Video

### v2

@@ Abstract
We introduce scenario-based cognitive status identification in older drivers from naturalistic driving videos, leveraging large vision models. In recent times, cognitive decline including Dementia and Mild Cognitive Impairment (MCI), is often underdiagnosed due to the time-consuming and costly nature of current diagnostic methods. By analyzing real-world driving behavior captured through in-vehicle sensors, this study aims to extract “digital fingerprints” that correlate with functional decline and clinical features of dementia. Moreover, modern large vision models can draw meaningful insights from everyday driving patterns across different roadway scenarios to early detect cognitive decline. We propose a framework that uses large vision models and naturalistic driving videos to analyze driver behavior, identify cognitive status and predict disease progression. We leverage the strong relationship between real-world driving behavior as an observation of the current cognitive status of the drivers where the vehicle can be utilized as a “diagnostic tool”. Our method identifies early warning signs of functional impairment, contributing to proactive intervention strategies. This work enhances early detection and supports the development of scalable, non-invasive monitoring systems to mitigate the growing societal and economic burden of cognitive decline in the aging population.
@@
Naturalistic driving, Vision model, Cognitive decline, Driving scenario
journal: anonymous
@@
[label4]organization=Department of Mechanical Engineering,addressline=Iowa State University, city=Ames, state=IA, country=USA
@@ 1 Introduction
Cognitive decline in the aging population is a critical health concern that manifests across a spectrum of stages. The pre-dementia stage, commonly referred to as mild cognitive impairment (MCI), is characterized by measurable cognitive impairment with minimal or no functional changes in daily living [1, 2]. Individuals at this stage (often in their 70s) may exhibit early signs of cognitive decline without severe disruptions to independence. As the condition progresses to dementia, more severe cognitive and functional impairments emerge, ultimately diminishing the ability to live independently [3]. While MCI is not exclusive to dementia, it often serves as an early indicator of the disease and highlights the need for early detection and intervention [4].
@@
The timely identification of MCI provides an opportunity to slow disease progression, improve quality of life, and reduce the societal burden associated with advanced stages of dementia. Traditional methods for diagnosing cognitive decline are often time-consuming, costly, and reliant on clinical visits or complex neuroimaging techniques, which may not be practical for widespread use [2, 5, 6]. To address this, our research explores an innovative approach using real-world driving behavior as a passive diagnostic tool [7, 8]. This study focuses on individuals exhibiting MCI due to mild dementia, as determined by clinical characteristics [9].
@@
Prior research on cognitive decline has explored a variety of methods to detect early changes in individuals at risk of dementia. Neuroimaging techniques, such as magnetic resonance imaging (MRI) [10, 11] and positron emission tomography (PET) [12, 13] have been widely used to identify structural and functional brain changes associated with dementia. These methods are often paired with biomarkers, such as amyloid-beta and tau proteins [14], blood-based biomarkers [15] to improve diagnostic accuracy. However, the high cost and limited accessibility of these techniques hinder their widespread application [16]. Wearable sensor technologies [17, 18] have also gained traction in recent years. Studies have demonstrated the utility of devices that monitor gait patterns, heart rate variability, and sleep quality to assess functional and cognitive impairments. Machine learning models [19, 20, 21] applied to sensor data have shown promise in predicting the progression of cognitive decline. However, these approaches often require participant compliance and are limited to specific contexts, such as controlled environments or clinical trials [22]. Some studies [23, 24, 25] demonstrated that everyday driving behavior, such as reduced driving distances, fewer unique destinations, and smaller driving spaces, is associated with symptomatic MCI. However, limited research has investigated the impact of preclinical MCI on driving behavior [26, 27], highlighting a critical gap in understanding early-stage cognitive decline. Recently, driving research has shifted towards naturalistic assessments using sensor devices installed in participants’ personal vehicles, enabling the collection of real-world driving behavior data [7, 26].
@@
While neuropsychological testing provides a comprehensive assessment of cognitive abilities, it often requires in-person clinical visits, participant compliance, and substantial time and resources — all of which may be challenging for individuals in the early stages of MCI [28]. In this context, our approach capitalizes on passive monitoring of real-world driving behavior, which offers a non-invasive, cost-effective, and scalable solution. Analyzing naturalistic driving data offers a promising alternative to capture subtle behavioral cues of cognitive decline in real-world settings [29, 30]. Studies show that behavioral patterns in everyday driving can be an alternative and effective approach to detect early-stage MCI, complementing traditional cognitive testing and potentially enabling earlier intervention [29, 31, 26]. By combining driving video data with advanced computer vision techniques, we aim to discover the relationship between driving behavior and cognitive assessment, providing a digital “diagnostic tool” for early detection of cognitive decline. We propose a scenario-based framework that identifies discriminative behavioral cues and shows how scenario-specific analysis can reveal early cognitive decline more effectively than aggregate driving metrics.
@@ 2.1.1 Participants
These data were derived from a cohort of community-dwelling older adults (aged 65 - 90 years) enrolled in longitudinal studies on dementia and driving behavior, conducted at the UNMC Mind and Brain Health Labs between March 2021 and January 2024. Recruitment procedures followed prior work [32, 29], involving community outreach efforts (e.g., flyers, social media, local news, discussions at senior organizations) and pre-screening to ensure eligibility based on defined inclusion criteria (see A). All participants lived independently in Nebraska or neighboring states (e.g., Iowa), held a valid driver’s license, were proficient in English and provided informed consent. Exclusion criteria were applied to minimize potential confounding effects from underlying health conditions and are detailed in A. The study protocol was approved by the UNMC Institutional Review Board (IRB 0522-20-FB) and all participants provided written informed consent prior to participation. A summary overview of the participants is shown in Table 1.
@@ 2.1.2 Data collection
Data collection includes recording driving video, sensor signals, actigraphy, sleep log, driving behavior and in-lab cognitive assessment conducted at UNMC. Over a 2-year period, driving data were collected in 3-month intervals using custom-installed “Black Box” systems in each participant’s personal vehicle. The system continuously recorded driving behavior throughout each trip over a 12-week period. It captured synchronized video (front roadway and cabin), vehicle sensor data (GPS, accelerometer, gyroscope), and on-board diagnostics data (speed, rpm, engine throttle). All data were recorded at a 1Hz sampling rate from engine start to stop (ignition cycle). Each ignition cycle indicates a distinct “drive” for the corresponding driver, excluding trips with zero mileage or speed<5<5mph (e.g., parking lot or driveway maneuvers) or if participants reported periods of non-driving (e.g., vacations, illness).
Table 1: Summary characteristics of the participants in RWRAD dataset
@@
Employment Employed full/part time: 19
@@
Driving Experience (years) Range: 45 - 78 58.7 6.6
@@ 2.1.3 Cognitive Status
Cognitive status was determined using the 2018 NIA-AA research criteria [9]. To assess cognitive status, participants completed a series of tests across five domains – memory, language, visuospatial abilities, executive function and attention (see B). Specialist clinicians in dementia and Alzheimer’s reviewed the neuropsychological evaluations for cognition and daily functioning from the National Alzheimer’s Coordinating Center (NACC) Unified Data Set (UDS) version 3 [33]. They reached a consensus on cognitive staging based on participants performance across the five cognitive domains. Raw test scores were converted into standardized $Z$-scores, adjusted for age, sex, and education. These $Z$-scores were then used to define cognitive impairment. As cognitive impairment progresses, it often becomes a multi-domain process, ultimately impacting functional abilities and transitioning to the dementia stage. The in-lab data collection focused on assessing functional abilities and activities through driving behavior in various roadway scenarios. Cognitive performance was evaluated using two composite cognitive scores. First, the COGSTAT score was utilized as a measure of cognitive function [34]. It was derived from 14 neuropsychological tests evenly distributed across those five cognitive domains. For each test, an age-adjusted $Z$-score was calculated and transformed into a $T$-score (mean = 50, SD = 10) using the formula $T=10(Z\times 10+50)$. The composite scores required complete data without missing tests to ensure validity. Second, the Montreal Cognitive Assessment (MoCA) score was used as a clinical screening tool designed to detect clinically significant cognitive change and age-related cognitive decline [35]. Both the MoCA and the full neuropsychological battery were administered to all participants. To address challenges associated with compliance and self-reporting accuracy in cognitively impaired participants, a study partner was recruited for each participant. After two weeks, participants returned to the study site with their partner for evaluations of cognitive status and daily functioning.
@@ 2.2 Video Data Processing
Figure 1: Breakdown of the driving data in the RWRAD dataset
Figure 2: Cognitive score (a) COGSTAT, (b) MoCA distribution across subjects ($N=69$). We combined the MD and MCI subjects into a unified class“MD-aging” and performed two-class classification with the “Normal-aging”.
Figure 3: GIS map of the Omaha most transited route database (Omaha, NE)
Figure 4: Scenario-specific video sampling. The top row represents the rollout of all valid drives and driving videos across subjects in the Omaha most transited route. From each subject $D_{i}$, valid drives $dr_{1},dr_{2},\dots,dr_{n}$} and their associated video samples $V_{1},V_{2},\dots,V_{m}$} were mapped into scenario-specific subsets. These video samples were aggregated and used as input to the large vision model for driving pattern analysis.
@@
Quality driving video is essential for reliable video-based modeling since missing or corrupted data can hurt the validity of findings. Ensuring rigorous processing of the collected video clips was therefore a crucial step before conducting our analysis. While processing the collected driving video data, we encountered some missing and corrupted video clips that required filtering and formatting. Figure 1 presents the breakdown of the processed driving data for the RWRAD dataset. The in-vehicle Blackbox system recorded a total of 520,662.7 minutes of driving data, segmented into valid drive duration (73.7%), Blackframe drive duration (3.9%) and Missing drive duration (22.4%). For the video-based analysis, only valid driving data was retained. Drives with missing or corrupted data and Black-frame data were excluded to ensure input data quality and consistency for driver behavior modeling.
@@
Subjects with their corresponding driving videos were filtered out for only those drives that occurred on the Omaha most transited route, a predefined set of road segments derived from the RWRAD dataset (see Figure 3). This criterion was applied to ensure adequate video data coverage for each participant and to make sure we have diverse driving scenario exposure in the video samples. In addition, the common driving route ensures that individual drivers commute on a shared set of road segments, reducing variability due to route differences. This allows the model to focus on relevant driving patterns rather than route features. For our analysis, a total of $N=69$ subjects were filtered out, after excluding the subjects without cognitive status labels and drivers outside the Omaha region. Figure 2 shows the distribution of the cognitive scores across three aging groups— MD, MCI, and Normal. Among 69 subjects, 35 drivers (26 MCI and 9 MD combined) were categorized as “MD-aging”, while 34 were classified as “Normal-aging”. This balanced distribution enabled a robust comparison between normal and MD-aging groups in terms of cognitive test scores and driving behavior. The MD-aging subjects consistently show the lowest cognitive scores, while normal-aging individuals cluster at the higher end. MCI subjects span the intermediate range, overlapping with both groups. It also highlights the transitional nature of cognitive decline and the inherent classification challenge.
@@
To ensure consistency and relevance in our analysis, we focused on short driving segments. We selected roadway segments where the drive distance can be $\leq 0.31$ mile or $\leq 500$ meters and defined it as a “route segment” (discussed in 3.1). These segments provide concentrated snapshots of driver behavior under controlled road conditions. This approach minimizes variability due to changes in driving context or prolonged periods of noisy observation. Additionally, the time of day and traffic conditions (peak vs. off-peak hours) were considered to account for environmental stressors that could uniquely challenge drivers with mild dementia. Next, the scenario-specific video sampling method is described in section 2.3.

### v3

@@ Abstract
We introduce scenario-based cognitive assessment in older drivers from naturalistic driving videos. In recent times, cognitive decline, including Dementia and Mild Cognitive Impairment (MCI), is often underdiagnosed due to the time-consuming and costly nature of current diagnostic methods. By analyzing real-world driving behavior captured through in-vehicle sensors, this study aims to extract “digital fingerprints” that reveal the correlation among driving behavior, functional decline and clinical features of dementia. Moreover, modern large vision models can draw meaningful insights from everyday driving patterns across different roadway scenarios. We propose a framework that utilizes large vision models and naturalistic driving videos captured in various driving scenarios to analyze driver behavior and assess cognitive status. We leverage the strong relationship between real-world driving behavior as an observation of the current cognitive state of the drivers, where the vehicle can be utilized as a “Diagnostic Tool”. The results indicate that our approach effectively captures the link between cognitive load and behavioral patterns. This work aims to develop a scalable and non-invasive monitoring strategy for detecting cognitive decline in the aging population.
@@
Naturalistic driving , Large Vision model , MCI , Dementia , Driving scenario
journal: a journal
@@
[label4]organization=Department of Mechanical Engineering,addressline=Iowa State University, city=Ames, state=IA, postcode=50011, country=USA
{graphicalabstract}
{highlights}
We introduce a novel Scenario-based Cognitive Assessment framework, utilizing naturalistic driving videos and large vision models to identify cognitive decline in older adults.
Our framework extracts discriminative visual patterns across diverse driving contexts and captures subtle behavioral cues to detect cognitive state.
This method demonstrates strong generalization and effectiveness in diverse driving scenarios and with unseen subjects, achieving up to 72.3% accuracy and an F1-score of 0.80 in the most contrastive scenario.
We highlight naturalistic driving behavior as a scalable and effective Diagnostic Tool which can offer passive cognitive health assessment for early dementia detection.
@@ 1 Introduction
Cognitive decline in the aging population is a critical health concern that manifests across a spectrum of stages. The pre-dementia stage, commonly referred to as Mild Cognitive Impairment (MCI), is characterized by measurable cognitive impairment with minimal or no functional changes in daily living [46, 61, 31]. Individuals often in their 70s may exhibit early signs of cognitive decline without severe disruptions to independence. As the condition progresses to dementia, more severe cognitive and functional impairments emerge, ultimately diminishing the ability to live independently [5, 67, 41]. While MCI is not exclusive to dementia, it often serves as an early indicator of the disease and highlights the need for early detection and intervention [14, 53].
@@
The timely identification of MCI provides an opportunity to slow disease progression, improve quality of life, and reduce the societal burden associated with advanced stages of dementia [44, 64, 49]. Traditional methods for diagnosing cognitive decline are often time-consuming, costly, and reliant on clinical visits or complex neuroimaging techniques, which may not be practical for widespread use [45, 61, 40, 4]. To address this, our research explores an innovative approach using real-world driving behavior as a passive diagnostic tool [6, 60]. This study focuses on individuals exhibiting MCI due to mild dementia, as determined by clinical characteristics [31].
@@
Prior research on cognitive decline has explored a variety of methods to detect early changes in dementia risk. Neuroimaging techniques, such as magnetic resonance imaging (MRI) [15, 33, 66], Electroencephalography (EEG) [3, 2] and positron emission tomography (PET) [57, 54] have been widely used to identify structural and functional brain changes associated with dementia. These methods are often paired with biomarkers, such as amyloid-beta and tau proteins [56, 34], blood-based biomarkers [24, 25] to improve diagnostic accuracy. However, the high cost and limited accessibility of these techniques hinder their widespread application [74]. Wearable sensors [62, 32, 69] have also gained traction in recent years. Studies have demonstrated the utility of devices that monitor gait patterns, heart rate variability, and sleep quality to assess functional and cognitive impairments. Machine learning models [72, 36, 42] applied to sensor data have shown promise in predicting the progression of cognitive decline. However, these approaches often require participant compliance and are limited to specific contexts, such as controlled environments or clinical trials [39, 63]. Some studies demonstrated that everyday driving behavior, such as reduced driving distances, fewer unique destinations, and smaller driving spaces, is associated with symptomatic MCI [68, 38, 17, 22]. However, limited research has investigated the impact of preclinical MCI on driving behavior [19, 55], highlighting a critical gap in understanding early-stage cognitive decline. Recently, driving research has shifted towards naturalistic assessments using sensor devices installed in participants’ personal vehicles, enabling the collection of real-world driving behavior data [6, 10, 13, 19].
While neuropsychological testing provides a comprehensive assessment of cognitive abilities, it often requires in-person clinical visits, participant compliance, and substantial time and resources — all of which may be challenging for individuals in the early stages of MCI [21, 47]. Recent naturalistic driving research, such as the LongROAD [19], demonstrated that combining driving variables with demographics can achieve up to 88% F1-score in predicting incident MCI or dementia. However, it still depends on aggregated trip-level statistics and lacks context-specific behavioral insights. Additionally, some recent studies [18, 23, 52, 68] have focused on scenario-dependent driving patterns analysis. For instance, cognitively impaired drivers tend to exhibit riskier behavior under adverse weather conditions [26]. These studies underscore the need for analyzing scenario-specific behaviors.
@@
Our method addresses these gaps by leveraging naturalistic video from specific driving contexts (e.g., freeway merging, stopping in a signalized intersection) and large vision models to extract nuanced, scenario-specific visual patterns. This setting offers detection of subtle behavioral cues revealed by the cognitive load of a driving scenario. Our approach highlights the importance of monitoring real-world driving behavior as a non-invasive, cost-effective, and scalable solution. Analyzing naturalistic driving data offers a promising alternative to capture subtle behavioral cues of cognitive decline in real-world settings [48, 28]. Studies show that behavioral patterns in everyday driving can be an alternative and effective approach to detect early-stage MCI, complementing traditional cognitive testing and potentially enabling earlier intervention [48, 20, 19]. By combining driving video data with advanced computer vision techniques, we aim to discover the relationship between driving behavior and cognitive assessment, providing a digital “diagnostic tool” for early detection of cognitive decline. We propose a scenario-based framework that identifies discriminative behavioral cues and shows how scenario-specific analysis can reveal early cognitive decline more effectively than aggregate driving metrics.
The main contributions of this paper are summarized as follows:
  • 1.
    To address the limitations of the traditional driving metrics-based approach in detecting early cognitive decline, we proposed a novel “Scenario-based Diagnostic Tool” that leverages naturalistic driving videos and a Large Vision model for cognitive status recognition in older adults.
  • 2.
    To analyze the driving contexts due to cognitive decline, we extracted discriminative visual features and leveraged a scenario-ranking strategy based on inter-group separability, which directly maps to the cognitive load of the driving task under stressful scenarios.
  • 3.
    To evaluate robustness and generalizability, we validated our framework under two sampling conditions - random sampling and driver-level separation (DLS) with cross-validation, showcasing consistent improvements in distinguishing cognitive aging groups across “unseen” participants.
  • 4.
    We highlighted the potential of scenario-based driving behavior monitoring as a scalable, cost-effective, and non-invasive “Diagnostic Tool” that offers clinical interventions for cognitive health assessment. Our framework achieved a maximum accuracy of 72.32% and a maximum F1-score of 0.80 (on the Freeway interchange scenario), outperforming a recent state-of-the-art approach on the same task. This result highlights the efficacy of scenario-based video analysis for early detection of cognitive decline.
The rest of this paper is organized as follows: Section 2 provides a comprehensive overview of the proposed method, including data processing, sampling protocols, vision model-based and scenario-based framework. Section 3 introduces the details of our experiment setup, including the large vision model backbone, scenario ranking strategy and classification model. Section 4 describes the experimental results and their implications for scenario-based cognitive aging-group classification. Finally, Section 5 concludes the paper.
@@ 2 Methodology
Figure 1: Scenario-based cognitive assessment framework. The pipeline consists of two phases - scenario selection and scenario validation. For testing, the fine-tuned RF model that produced the highest performance was used.
We proposed a scenario-based framework for cognitive status assessment from naturalistic driving videos. As illustrated in Fig. 1, the pipeline consists of five modules. First, (1) raw driving videos were collected, filtered, and aligned to the most-transited routes (discussed in section 2.2) to ensure consistent scenario coverage across subjects. Next, (2) a pretrained large vision model encodes each video segment to vectors that were further projected to low-dimensional feature representations ($Z_{k}$). Then, (3) scenarios were ranked based on the Euclidean distance between the Normal and MD-aging groups in the low-dimensional space. The scenario with the largest Euclidean distance (most contrastive) and the smallest distance (least contrastive) were selected. To evaluate the ranking, (4) the most and least contrastive scenario vectors were then used to train two identical Random Forest (RF) models [9]. Finally, (5) performance difference between these RF models was observed as they indicate the validity of the scenario-ranking and contrastive scenario selection step. A large performance gap indicates most contrastive scenario pair. During testing, the best-performing RF model was utilized to classify the cognitive groups. In this section, we will discuss each module and model configuration in detail.
@@ 2.1.1 Participants
These data were derived from a cohort of community-dwelling older adults (aged 65 - 90 years) enrolled in longitudinal studies on dementia and driving behavior, conducted at the UNMC Mind and Brain Health Labs between March 2021 and January 2024. Recruitment procedures followed prior work [12, 48], involving community outreach efforts (e.g., flyers, social media, local news, discussions at senior organizations) and pre-screening to ensure eligibility based on defined inclusion criteria (see section A). All participants lived independently in Nebraska or neighboring states (e.g., Iowa), held a valid driver’s license, were proficient in English and provided informed consent. Exclusion criteria were applied to minimize potential confounding effects from underlying health conditions and are detailed in section A. The study protocol was approved by the UNMC Institutional Review Board (IRB 0522-20-FB) and all participants provided written informed consent prior to participation. A summary overview of the participants is shown in Table 1.
@@ 2.1.2 Data collection
Data collection includes recording driving video, sensor signals, actigraphy, sleep log, driving behavior and in-lab cognitive assessments conducted at UNMC. Over a two-year period, driving data were collected in three-month intervals using custom-installed “Black Box” systems in each participant’s personal vehicle. The system continuously recorded driving behavior throughout each trip over a 12-week period. It captured synchronized video (front roadway and cabin), vehicle sensor data (GPS, accelerometer, gyroscope), and on-board diagnostics data (speed, rpm, engine throttle). All data were recorded at a 1 Hz sampling rate from engine start to stop (ignition cycle). Each ignition cycle indicates a distinct “drive” for the corresponding driver, excluding trips with zero mileage or speed <5<5mph (e.g., parking lot or driveway maneuvers) or if participants reported periods of non-driving (e.g., vacations, illness).
Table 1: Summary characteristics of the participants in the RWRAD dataset
@@
Employment Employed full/part-time: 19
@@
Driving Exp. (years) Range: 45 - 78 58.7 6.6
@@ 2.1.3 Cognitive Status
Cognitive status was determined using the 2018 NIA-AA research criteria [31]. To assess cognitive status, participants completed a series of tests across five domains – memory, language, visuospatial abilities, executive function and attention (see section B). Specialist clinicians in dementia and Alzheimer’s reviewed the neuropsychological evaluations for cognition and daily functioning from the National Alzheimer’s Coordinating Center (NACC) Unified Data Set (UDS) version 3 [1]. They reached a consensus on cognitive staging based on participants performance across the five cognitive domains. Raw test scores were converted into standardized $Z$-scores, adjusted for age, sex, and education. These $Z$-scores were then used to define cognitive impairment. As cognitive impairment progresses, it often becomes a multi-domain process, ultimately impacting functional abilities and transitioning to the dementia stage. The in-lab data collection focused on assessing functional abilities and activities through driving behavior in various roadway scenarios. Cognitive performance was evaluated using two composite cognitive scores. First, the COGSTAT score was utilized as a measure of cognitive function [59]. It was derived from 14 neuropsychological tests evenly distributed across those five cognitive domains. For each test, an age-adjusted $Z$-score was calculated and transformed into a $T$-score (mean = 50, SD = 10) using the formula $T=10(Z\times 10+50)$. The composite scores required complete data without missing tests to ensure validity. Second, the Montreal Cognitive Assessment (MoCA) score was used as a clinical screening tool designed to detect clinically significant cognitive change and age-related cognitive decline [51]. Both the MoCA and the full neuropsychological battery were administered to all participants. To address challenges associated with compliance and self-reporting accuracy in cognitively impaired participants, a study partner was recruited for each participant. After two weeks, participants returned to the study site with their partner for evaluations of cognitive status and daily functioning.
@@ 2.2 Video Data Processing
Figure 2: Breakdown of the recorded driving data in the RWRAD dataset.
Figure 3: Cognitive score distribution across all subjects ($N=69$). (a) COGSTAT score, (b) MoCA score. We combined the MD and MCI subjects into a unified class “MD-aging” and performed two-class classification with the “Normal-aging”.
Figure 4: Geospatial map of the OMR database (Omaha, NE). The purple lines indicate the routes where the participants drove most frequently during the RWRAD data collection period.
Figure 5: Scenario-specific video sampling. The top row represents the rollout of all valid drives and driving videos across subjects in the OMR data. From each subject $D_{i}$, valid drives $dr_{1},dr_{2},\dots,dr_{n}$} and their associated video samples $V_{1},V_{2},\dots,V_{m}$} were mapped into scenario-specific subsets. These video samples were aggregated and used as input to the large vision model for driving pattern analysis.
@@
Quality driving video is essential for reliable video-based modeling since missing or corrupted data can hurt the validity of findings. Ensuring rigorous processing of the collected video clips was therefore a crucial step before conducting our analysis. While processing the collected driving video data, we encountered some missing and corrupted video clips that required filtering and formatting. Fig. 2 presents the breakdown of the processed driving data for the RWRAD dataset. The in-vehicle Blackbox system recorded a total of 520,662.7 minutes of driving data, segmented into valid drive duration (73.7%), Blackframe drive duration (3.9%) and Missing drive duration (22.4%). For the video-based analysis, only valid driving data was retained. Drives with missing or corrupted data and Black-frame data were excluded to ensure quality and consistency of the input data.
@@
Subjects with their corresponding driving videos were further filtered out for only those drives that occurred on the Omaha Most-transited Route (OMR), a predefined set of road segments derived from the RWRAD dataset (shown in Fig. 4). This criterion was applied to ensure adequate video data coverage for each participant and to make sure we have diverse driving scenario exposure in the video samples. In addition, the common driving route ensures that individual drivers commute on a shared set of road segments, reducing variability due to route differences. This allows the model to focus on relevant driving patterns rather than route features. For our analysis, a total of $N=69$ subjects were filtered out, after excluding the subjects without cognitive status labels and drivers outside the Omaha region. Fig. 3 shows the distribution of the cognitive scores across three aging groups— MD, MCI, and Normal. Among 69 subjects, 35 drivers (26 MCI and 9 MD combined) were categorized as “MD-aging”, while 34 were classified as “Normal-aging”. This balanced distribution enabled a robust comparison between normal and MD-aging groups in terms of cognitive test scores and driving behavior. The MD-aging subjects consistently show the lowest cognitive scores, while normal-aging individuals cluster at the higher end. MCI subjects span the intermediate range, overlapping with both groups. It also highlights the transitional nature of cognitive decline and the inherent challenge in classification.
@@
To ensure consistency and relevance in our analysis, we focused on short driving segments. We selected roadway segments where the drive distance is $\leq 0.31$ mile or $\leq 500$ meters and defined them as “route segment” (discussed in section 3.1). These segments provide concentrated snapshots of driver behavior under controlled road conditions. This approach minimizes variability due to changes in driving context or prolonged periods of noisy observation. Additionally, the time of day and traffic conditions (peak vs. off-peak hours) were considered to account for environmental stressors that could uniquely challenge drivers with mild dementia. Next, the scenario-specific video sampling method is described in section 2.3.